find mean, variance, standard deviation of n= 120 and p= 0.4
also can u explain how you get that little extra number in the variance, im not understanding i get stuck there.

Answer by stanbon(75887)

find mean, variance, standard deviation of
n= 120 and p= 0.4 and q = 0.6
---
mean = np = 0.4*120 = 48
---
std = sqrt(npq) = sqrt(48*0.6) = 5.3666..
---
variance = npq = 48*0.6 = 28.8
==================================
Cheers,
Stan H.
===============